# AMBER: An Adaptive Multimodal Mask Transformer for Beam Prediction with Missing Modalities

### v1

@@ I Introduction
To improve robustness, multi-modal fusion has emerged as a powerful technique for robust beam prediction by combining complementary information from multiple sensors such as cameras, LiDAR, radar, and GPS [16, 17, 18, 19, 20]. For instance, [16] extracted visual features via a residual network (ResNet) and fused them with positional information through a multilayer perceptron (MLP) for beam prediction, demonstrating improved performance compared to single-modality methods. The cross-modal feature enhancement framework in [17] instead jointly processed image and radar inputs using an uncertainty-aware dynamic fusion module and leveraged temporal correlations across multiple frames, which resulted in superior performance to earlier fusion approaches. Deeper multimodal integration has since been pursued: [19] combined feature vectors extracted from radar, camera, LiDAR, and GPS data, with historical beam indices, and learned their temporal dependencies using an long short-term memory (LSTM) network; [21] proposed a hybrid deep quantum-transformer network (QTN) for beam prediction; [18] proposed a multimodal transformer architecture to fuse intermediate-scale features from multiple sensors to better capture intermodal relations and temporal patterns; and [22] introduced a lightweight multimodal fusion framework that exploits the linear complexity of state space models. Although effective, all these works assumed full modality availability, an assumption often violated in practice due to sensor blockage, adverse lighting, weather conditions, or GPS outages. Consequently, their performance degrades severely when one or more modalities are missing.
@@
Figure 1: An illustration of the system model, where the BS is equipped with a mmWave antenna array and also with a camera, a LiDAR, and a Radar, while the single-antenna UE is equipped with a GPS.
@@
Notations: Scalars, vectors, and matrices are denoted by ower/upper case, bold-face lower-case, and bold-face upper-case letters, respectively. The superscript $(\cdot)^{H}$ and $(\cdot)^{T}$ represent the Hermitian and tranpose operations, respectively. The symbol $|\cdot|$ represents the magnitude of a scalar. The symbol $\odot$ denotes element-wise multiplication. The indicator function $\bf{1}\{\cdot\}$ returns 1 when its condition is satisfied, and 0 otherwise. The sets $\mathbb{R}$ and $\mathbb{C}$ denote the real and complex numbers, respectively. Operators $\text{min}(\cdot)$ and $\text{max}(\cdot)$ return the smallest and largest values of their arguments, respectively. $\text{FFT}_{\text{2D}}$ denotes the 2D-FFT operation. The squared $L_{2}$ norm of a vector $\bf{x}$} is written as $\|{\bf{x}}\|_{2}^{2}=\sum_{i}x_{i}^{2}$. The operator $\text{Concat}(\cdot)$ defines tokens concatenating along the feature dimension, forming a single composite input token. The sigmoid function is defined as $\text{Sig}(x)=\frac{1}{1+e^{-x}}$. The softmax function maps a vector $\bf{x}$} to a probability distribution according to $\text{softmax}({\bf{x}})_{i}=\frac{\text{exp}({{x}_{i}})}{\sum_{i}\text{exp}({{x}}_{i})}$, where $(\cdot)_{i}$ denotes the $i$-th element. The operator $\text{exp}(\cdot)$ denotes the exponential function with base $e$, i.e., $\text{exp}(x)=e^{x}$.
@@ II-A System Model
  • ${\bf{X}}_{\text{P}}[t]\in\mathbb{R}^{2\times 1}$: GPS data, providing the latitude and longitude coordinates of the UE.
@@ II-B Problem Formulation
Traditional pilot-based beam training can achieve this objective, but it becomes prohibitively costly for large codebooks with narrow beams, as a substantial number of pilots are required. Since the optimal beam direction strongly depends on the geometric relationship among the BS, UE, and surrounding environment, DL-based beam prediction offers a promising alternative by levering multimodal sensing information. The BS collects heterogeneous sensor data, including camera images, LiDAR point clouds, radar reflections, and positional coordinates as mentioned above, to provide rich contextual cues for learning environment-beam relationships. However, practical factors such as sensor blockage, adverse lighting, or weather conditions may render some modalities unavailable. To model this, let $\mathcal{I}\in\{\text{img, Lidar, radar, beam, pos}\}$ be the set of sensing and communication modalities. For each modality $i\in\mathcal{I}$, we define an availability indicator $m_{i}\in\{0,1\}$, where $m_{i}=1$ indicates that the corresponding modality is available and $m_{i}=0$ otherwise. The modality-availability vector is then expressed as
${\bf{m}}=[m_{\text{img}},m_{\text{Lidar}},m_{\text{radar}},m_{\text{beam}},m_{\text{pos}}]\in\{0,1\}^{5}.$ (3)
To capture both spatial and temporal correlations across different time instances, the BS collects historical multimodal observations $\{{\bf{X}}_{I}[t_{1}],{\bf{X}}_{L}[t_{1}],{\bf{X}}_{R}[t_{1}],{\bf{X}}_{P}[t_{2}]\}$, where $t_{1}\in\{t-W+1,t-W+2,\ldots,t\}$ span a sliding window of size $W$, and $t_{2}\in\{t-1,t\}$. In addition, historical beam indices ${\bf{X}}_{B}[t_{3}]$ with $t_{3}\in\{t-W,t-W+1,\ldots,t-1\}$, also serve as side information and are readily available at the BS to facilitate beam prediction.
@@
${\hat{k}^{*}[t]}={{\mathcal{P}}}({\bf{X}}_{\text{I}}[t_{1}],{\bf{X}}_{\text{L}}[t_{1}],{\bf{X}}_{\text{R}}[t_{1}],{\bf{X}}_{\text{P}}[t_{2}],{\bf{X}}_{\text{B}}[t_{3}],{\bf{m}}[t_{1}];\bf{\Theta}).$ (4)
Figure 2: An illustration of the proposed AMBER: The BS collects multimodal data, including images, LiDAR, radar, and GPS, along with historical beam indices. After modality-specific encoders extract initial features, time-aware positional embeddings are incorporated to preserve temporal dependencies. The augmented features are then processed by an adaptive multimodal mask transformer, which handles arbitrary missing modalities, performs cross-modal fusion and processs by CMA module. Finally, the resulting fusion token is used to predict the optimal beam index.
@@ III-A Modality-Specific Encoders
Radar Encoder: For radar data processing, we adopt the method in [25] to extract the range, angle and velocity information of the UE. Specifically, range–angle (RA) and range–velocity (RV) maps are generated using two-dimensional fast Fourier transform (2D-FFT) as follows:
${\bf{X}}_{\text{R}}^{RA}[t]=\sum_{a=1}^{A_{R}}\left|\text{FFT}_{\text{2D}}({\bf{X}}_{\text{R}}[t][:,:,a])\right|,$
@@
To preserve both spatial and velocity features, the RA and RV maps are concatenated along the channel dimension, yielding
$\bar{\bf{X}}_{\text{R}}[t]$ $=\text{Concat}\left({\bf{X}}_{\text{R}}^{RA}[t],{\bf{X}}_{\text{R}}^{RV}[t]\right).$ (6)
After min-max normalization, the preprocessed data are fed into a ResNet18-based encoder ${\mathcal{P}}_{\text{R}}$} [26] (with classification layers removed), followed by an adaptive pooling operator ${{{\mathcal{P}}_{\text{ada}}}}:\mathbb{R}^{C\times H_{\text{in}}\times W_{\text{in}}}\rightarrow\mathbb{R}^{C\times V_{A}\times H_{A}}$, which adjusts the pooling kernel size to $\lceil\frac{H_{\text{in}}}{V_{A}}\rceil\times\lceil\frac{W_{\text{in}}}{H_{A}}\rceil$ to maintain feature topology consistency. Here, $V_{A}$ and $H_{A}$ represent the predefined encoded feature dimensions, respectively. The radar encoder input-output is therefore defined as
${\bf{x}}_{\text{R}}[t]={{\mathcal{P}}_{\text{ada}}}\left({{\mathcal{P}}_{\text{R}}}(\bar{\bf{X}}_{\text{R}}[t])\right)\in\mathbb{R}^{C\times V_{A}\times H_{A}}.$ (7)
@@
LiDAR Encoder: LiDAR data ${\bf{X}}_{L}[t]$ consists of raw 3D point clouds providing precise spatial geometry information. To obtain an efficient representation, the point cloud is projected onto a bird’s-eye-view (BEV) grid of size $V_{g}\times H_{g}$ to form a histogram ${\bf{X}}^{\text{H}}_{L}[t]$, where the maximum number of points per grid cell is capped at five to reduce outlier effects, and $V_{g}$ and $H_{g}$ are predefined constants to control grid size. The BEV representation preserves the spatial layout of the environment in a top-down perspective while compressing the 3D geometry into a computationally tractable 2D format. The histogram is normalized and then fed into a modified ResNet18 backbone with adaptive pooling, defined as
${\bf{x}}_{\text{L}}[t]={\mathcal{P}_{\text{L}}}\left(\mathcal{N}\left({\bf{X}}^{\text{H}}_{L}[t]\right)\right),$ (9)
where $\mathcal{N}(\cdot)$ denotes the min–max normalization operator and is calculated as
@@ III-B1 Input Embeddings
To capture the varying contribution of different modalities and handle missing-modality cases, we further introduce a feature reweighting indicator ${\mathcal{P}}_{\text{ind}}$ to construct reweighted modality tokens as
${\bm{\xi}}=\text{Sig}({\mathcal{P}}_{\text{ind}}({\bm{\xi}}))\odot{\bm{\xi}},$ (17)
which is an activation function that maps the input to a value between 0 and 1. To avoid over-reliance on any single modality, particularly under random modality dropout, an $L_{2}$ regularization term is added to the loss function (formally introduced later in (38)) to promote balanced feature utilization, which is formulated as
$L_{2}=\|\text{Sig}({\mathcal{P}}_{\text{ind}}({\bm{\xi}}))\|_{2}^{2}.$ (18)
@@ III-B2 Modality-specific Block
To facilitate efficient parallel computation arcoss all modalities, we concatenate their input tokens into a unified sequence $\bm{\xi}$. Accordingly, a binary mask matrix ${\bf{M}}_{j,i}$ is designed based on $\bf{m}$ in (3) to prevent noise propagation from missing modalities, which is illustrated in Fig. 3. Specifically, ${\bf{M}}_{j,i}=1$ if modality $j$ can attend to modality $i$, and ${\bf{M}}_{j,i}=0$ otherwise. Then, we have
@@
$\text{head}_{j,h}=\sum_{i}\left(\frac{e^{\frac{{\bf{Q}}_{j,h}{\bf{K}}_{i,h}^{T}}{\sqrt{d_{k}}}}{\odot{\bf{M}}_{j,i}}}{\sum_{i^{\prime}}e^{\frac{{\bf{Q}}_{j,h}{\bf{K}}^{T}_{i^{\prime},h}}{\sqrt{d_{k}}}}\odot{\bf{M}}_{j,{i^{\prime}}}}\right){\bf{V}}_{i}.$ (24)
@@ III-B3 Modality-fusion Block
${\bf{Z}}^{\prime}_{F}=\text{LN}(\text{MCA}([{\bm{\xi}}_{F},{\bm{\xi}}])+{\bm{\xi}}_{F}).$ (29)

### v2

@@ I Introduction
To improve robustness, multi-modal fusion has emerged as a powerful technique for robust beam prediction by combining complementary information from multiple sensors such as cameras, LiDAR, radar, and GPS [16, 17, 18, 19, 20, 21, 22, 23]. For instance, [18] extracted visual features via a residual network (ResNet) and fused them with positional information through a multilayer perceptron (MLP) for beam prediction, demonstrating improved performance compared to single-modality methods. The cross-modal feature enhancement framework in [19] instead jointly processed image and radar inputs using an uncertainty-aware dynamic fusion module and leveraged temporal correlations across multiple frames, which resulted in superior performance to earlier fusion approaches. Deeper multimodal integration has since been pursued: [21] combined feature vectors extracted from radar, camera, LiDAR, and GPS data, with historical beam indices, and learned their temporal dependencies using an long short-term memory (LSTM) network; [24] proposed a hybrid deep quantum-transformer network (QTN) for beam prediction; [20] proposed a multimodal transformer architecture to fuse intermediate-scale features from multiple sensors to better capture intermodal relations and temporal patterns; [25] introduced a lightweight multimodal fusion framework that exploits the linear complexity of state space models; and [23] introduced ENWAR 2.0 through task-oriented explainability metrics, showing that its adaptive RAG-enabled LLM framework improves beam interpretation and environmental perception in terms of faithfulness, correctness, and answer relevancy. Although effective, all these works assumed full modality availability, an assumption often violated in practice due to sensor blockage, adverse lighting, weather conditions, or GPS outages. Consequently, their performance degrades severely when one or more modalities are missing.
@@
Figure 1: An illustration of the system model.
@@
Notations: Scalars, vectors, and matrices are denoted by ower/upper case, bold-face lower-case, and bold-face upper-case letters, respectively. The superscript $(\cdot)^{H}$ and $(\cdot)^{T}$ represent the Hermitian and transpose operations, respectively. The symbol $|\cdot|$ represents the magnitude of a scalar. The symbol $\odot$ denotes element-wise multiplication. The indicator function $\bf{1}\{\cdot\}$ returns 1 when its condition is satisfied, and 0 otherwise. The sets $\mathbb{R}$ and $\mathbb{C}$ denote the real and complex numbers, respectively. Operators $\text{min}(\cdot)$ and $\text{max}(\cdot)$ return the smallest and largest values of their arguments, respectively. $\text{FFT}_{\text{2D}}$ denotes the 2D-FFT operation. The squared $L_{2}$ norm of a vector $\bf{x}$} is written as $\|{\bf{x}}\|_{2}^{2}=\sum_{i}x_{i}^{2}$. The operator $\text{Concat}(\cdot)$ defines tokens concatenating along the feature dimension, forming a single composite input token. The sigmoid function is defined as $\text{Sig}(x)=\frac{1}{1+e^{-x}}$. The softmax function maps a vector $\bf{x}$} to a probability distribution according to $\text{softmax}({\bf{x}})_{i}=\frac{\text{exp}({{x}_{i}})}{\sum_{i}\text{exp}({{x}}_{i})}$, where $(\cdot)_{i}$ denotes the $i$-th element. The operator $\text{exp}(\cdot)$ denotes the exponential function with base $e$, i.e., $\text{exp}(x)=e^{x}$.
@@ II-A System Model
  • ${\bf{X}}_{\text{G}}[t]\in\mathbb{R}^{2\times 1}$: GPS data, providing the latitude and longitude coordinates of the UE.
@@ II-B Problem Formulation
Traditional pilot-based beam training can achieve this objective, but it becomes prohibitively costly for large codebooks with narrow beams, as a substantial number of pilots are required. Since the optimal beam direction strongly depends on the geometric relationship among the BS, UE, and surrounding environment, DL-based beam prediction offers a promising alternative by levering multimodal sensing information. The BS collects heterogeneous sensor data, including camera images, LiDAR point clouds, radar reflections, and positional coordinates as mentioned above, to provide rich contextual cues for learning environment-beam relationships. However, practical factors such as sensor blockage, adverse lighting, or weather conditions may render some modalities unavailable. To model this, let $\mathcal{I}=\{I,L,R,B,G\}$ denote the set of sensing and communication modalities, corresponding to image, LiDAR, radar, beam, and GPS, respectively, with cardinality number I. For each modality $i\in\mathcal{I}$, we define an availability indicator $m_{i}\in\{0,1\}$, where $m_{i}=1$ indicates that the corresponding modality is available and $m_{i}=0$ otherwise. The modality-availability vector is then expressed as
${\bf m}=[m_{I},m_{L},m_{R},m_{B},m_{G}]\in\{0,1\}^{5}.$ (3)
To capture both spatial and temporal correlations, the BS collects historical multimodal observations over a sliding window of length $W$. For notational simplicity, we define $\mathbf{X}_{j}^{W}[t]\triangleq\{X_{j}[\tau]\}_{\tau=t-W+1}^{t}$ for $j\in\{I,L,R\}$ and $\mathbf{X}_{G}^{W}[t]\triangleq\{X_{G}[\tau]\}_{\tau=t-1}^{t}$. In addition, historical beam indices $\mathbf{X}_{B}^{W}[t-1]\triangleq\{X_{B}[\tau]\}_{\tau=t-W+1}^{t-1}$ also serve as side information and are readily available at the BS to facilitate beam prediction.
@@
$\hat{k}^{*}[t]\!=\!\mathcal{P}\!\left(\mathbf{X}_{I}^{W}[t],\mathbf{X}_{L}^{W}[t],\mathbf{X}_{R}^{W}[t],\mathbf{X}_{G}^{W}[t],\mathbf{X}_{B}^{W}[t-1],\mathbf{m}[t];\bf{\Theta}\!\right).$ (4)
@@ III Robust Multimodal Beam Prediction With Missing Modalities
Figure 2: Illustration of the proposed AMBER: The BS collects multimodal data, including images, LiDAR, radar, and GPS, along with historical beam indices. Modality-specific encoders are first employed to extract initial feature representations. Time-aware positional embeddings are then incorporated to capture temporal dependencies. The resulting features are fed into an adaptive multimodal mask transformer, which accommodates arbitrary missing modalities and extracts a unified fusion representation. Finally, the resulting fusion token is used to predict the optimal beam index.
@@ III-A Modality-Specific Encoders
Radar Encoder: For radar data processing, we adopt the method in [28] to extract the range, angle and velocity information of the UE from the raw radar tensor ${\bf{X}}_{\text{R}}[t]\in\mathbb{R}^{M_{R}\times S_{R}\times A_{R}}$. Specifically, range–angle (RA) and range–velocity (RV) maps are generated using two-dimensional fast Fourier transform (2D-FFT). To ensure consistent spatial dimensions, the angle and velocity axes are zero-padded to a fixed size $N_{\text{FFT}}$ prior to performing the FFT. The RA and RV maps are computed as
${\bf{X}}_{\text{R}}^{RA}[t]$ $=\left(\sum_{a=1}^{A_{R}}\left|\text{FFT}_{\text{2D}}({\bf{X}}_{\text{R}}[t][:,:,a])\right|\right)^{T},$
@@
Both ${\bf{X}}_{\text{R}}^{RA}[t]$ and ${\bf{X}}_{\text{R}}^{RV}[t]$ share the same spatial dimensions $S_{R}\times N_{\text{FFT}}$. To jointly preserve spatial and velocity features without information loss, these two 2D maps are treated as single-channel images by introducing an explicit channel dimension (i.e., reshaped to $1\times S_{R}\times N_{\text{FFT}}$). They are then concatenated along the channel dimension to form
$\bar{\bf{X}}_{\text{R}}[t]$ $=\text{Concat}\left({\bf{X}}_{\text{R}}^{RA}[t],{\bf{X}}_{\text{R}}^{RV}[t]\right)\in\mathbb{R}^{2\times S_{R}\times N_{\text{FFT}}}.$ (6)
Next, min-max normalization is applied:
$\tilde{\bf{X}}_{\text{R}}[t]=\mathcal{N}(\bar{\bf{X}}_{\text{R}}[t]),$ (7)
where $\mathcal{N}(\cdot)$ denotes the normalization operator defined as
The normalized tensor $\tilde{\bf{X}}_{\text{R}}[t]$ is then fed into a ResNet18-based encoder ${\mathcal{P}}_{\text{R}}$} [29], whose classification head is removed and whose first convolutional layer is modified to accept the above two-channel radar input instead of the standard three-channel RGB input. Let ${{\mathcal{P}}_{\text{R}}}(\tilde{\bf{X}}_{\text{R}}[t])\in\mathbb{R}^{C\times H_{\mathrm{in}}\times W_{\mathrm{in}}}$ denote the output feature of the encoder, where $C$, $H_{\mathrm{in}}$, and $W_{\mathrm{in}}$ represent the number of channels, height, and width, respectively. To obtain a fixed-size representation, an adaptive pooling operator ${\mathcal{P}}_{\mathrm{ada}}:\mathbb{R}^{C\times H_{\mathrm{in}}\times W_{\mathrm{in}}}\rightarrow\mathbb{R}^{C\times V_{A}\times H_{A}}$ is then applied, where $V_{A}$ and $H_{A}$ represent the predefined output dimensions. The final radar encoder output is given by
${\bf{x}}_{\text{R}}[t]={{\mathcal{P}}_{\text{ada}}}\left({{\mathcal{P}}_{\text{R}}}(\tilde{\bf{X}}_{\text{R}}[t])\right).$ (9)
@@
LiDAR Encoder: LiDAR data ${\bf{X}}_{L}[t]$ consists of raw 3D point clouds providing precise spatial geometry information. To obtain an efficient representation, the point cloud is projected onto a bird’s-eye-view (BEV) grid of size $V_{g}\times H_{g}$ to form a histogram ${\bf{X}}^{\text{H}}_{L}[t]$, where the maximum number of points per grid cell is capped at five to reduce outlier effects, and $V_{g}$ and $H_{g}$ are predefined constants to control grid size. The BEV representation preserves the spatial layout of the environment in a top-down perspective while compressing the 3D geometry into a computationally tractable 2D format. The histogram is normalized and then fed into a modified ResNet18 backbone, where the first convolutional layer of ResNet18 is modified to accept input dimension, while the remaining residual architecture is kept unchanged and the classification head is removed. After adaptive pooling, the output is defined as
${\bf{x}}_{\text{L}}[t]={\mathcal{P}_{\text{L}}}\left(\mathcal{N}\left({\bf{X}}^{\text{H}}_{L}[t]\right)\right).$ (11)
@@ III-B1 Input Embeddings
To capture the varying contribution of different modalities and handle missing-modality cases, we further introduce a learnable modality-weight indicator ${\mathcal{P}}_{\text{ind}}$. Specifically, ${\mathcal{P}}_{\text{ind}}$ is parameterized by a trainable weight vector $\mathbf{w}_{\text{mod}}=[w_{I},w_{L},w_{R},w_{B},w_{G}]$, this vector is normalized through a temperature-controlled softmax operation:
$\bm{\alpha}={\mathcal{P}}_{\text{ind}}(\mathbf{w}_{\text{mod}})=\text{Softmax}\!\left(\mathbf{w}_{\text{mod}}/\tau_{\text{mod}}\right).$ (18)
where $\tau_{\text{mod}}$ is the temperature hyperparameter. The normalized weights satisfy $\alpha_{i}\geq 0$ and $\sum_{i\in\mathcal{I}}\alpha_{i}=1$. Each modality token is then reweighted as
${\bm{\xi}}_{i}=\alpha_{i}\bm{\xi}_{i},\quad i\in\mathcal{I}.$ (19)
To avoid over-reliance on any single modality, particularly under random modality dropout, an $L_{2}$ regularization term is added to the loss function (formally introduced later in (40)) to promote balanced feature utilization. By applying the modality-availability vector $\bf{m}$ to filter out missing modalities, this regularization loss is formulated as
$L_{2}=\sum_{i\in\mathcal{I}}m_{i}\,w_{i}^{2},$ (20)
@@ III-B2 Modality-specific Block
To facilitate efficient parallel computation across all modalities, we concatenate their input tokens into a unified sequence $\bm{\xi}$. Accordingly, a binary mask matrix ${\bf{M}}_{j,i}$ is designed based on $\bf{m}$ in (3) to prevent noise propagation from missing modalities, which is illustrated in Fig. 3. Specifically, ${\bf{M}}_{j,i}=1$ if modality $j$ can attend to modality $i$, and ${\bf{M}}_{j,i}=0$ otherwise. Then, we have
@@
$\text{head}_{j,h}=\sum_{i}\left(\frac{e^{\frac{{\bf{Q}}_{j,h}{\bf{K}}_{i,h}^{T}}{\sqrt{d_{k}}}}{\odot{\bf{M}}_{j,i}}}{\sum_{i^{\prime}}e^{\frac{{\bf{Q}}_{j,h}{\bf{K}}^{T}_{i^{\prime},h}}{\sqrt{d_{k}}}}\odot{\bf{M}}_{j,{i^{\prime}}}}\right){\bf{V}}_{i,h}.$ (26)
@@ III-B3 Modality-fusion Block
${\bf{Z}}^{\prime}_{F}=\text{LN}(\text{MCA}({\bm{\xi}}_{F},{\bm{\xi}})+{\bm{\xi}}_{F}).$ (31)
@@ III-B4 Class-former-aided Modality Alignment
where $\mathrm{CA}(\cdot)$ denotes a standard cross-attention operation, the first input is the query token and the second input provides the key and value tokens.
@@
III-C Beam Prediction
[…]
The overall training loss function is defined as
$L=\lambda_{f}L_{f}+\lambda_{c}L_{c}+\lambda_{r}L_{2},$ (40)
@@ III-C Beam Prediction
IV Simulation Results
IV-A Simulation Settings
IV-A1 Dataset and training setup
To evaluate the proposed AMBER framework under realistic mmWave communication conditions, we conduct experiments using the real-world DeepSense 6G dataset [27]. To the best of our knowledge, this dataset is the first large-scale real-world sensing-communication dataset that provides synchronized multimodal sensing and communication measurements. Owing to its realism and comprehensive data modalities, it has been widely adopted as a representative benchmark for sensing-aided beam prediction tasks.